OpenStudy (anonymous):

If (2x-5)/(x-4) = 3, then x = 7 What is wrong with this algebraic proof? suppose x = 7, then (2*7-5)/(7-4) = 3. Therefor, if (2x-5)/(x-4) = 3, then x = 7.

OpenStudy (anonymous):

I know you can't assume the conclusion, but isn't what we do to check our solutions?

hartnn (hartnn):

thats trial and error method

hartnn (hartnn):

yes, thats how we make sure that x=7 is a solution.

OpenStudy (anonymous):

I agree with @hartnn

Parth (parthkohli):

The thing that is wrong with this proof is that it uses the converse of a statement, which is not true in all cases. Let me give you an example: If \(x = 3\), \(x^2 - 10x + 21 = 0\). This one is true. However, If \(x^2 - 10x + 21 = 0\), \(x = 3\) is not true at all.

hartnn (hartnn):

its partly true x=3 or x=7 is completely true

Parth (parthkohli):

Hehe, but it's a false statement, as there are only true and false here.

hartnn (hartnn):

i would have gone with true, because its an "or" there and not "and"

hartnn (hartnn):

If x2−10x+21=0, x=3 TRUE

Parth (parthkohli):

No, not talking about your statement. I'm talking about mine... what I meant was "partly true" isn't a thing.

OpenStudy (anonymous):

huhm... a statement can't be partially true though

hartnn (hartnn):

what i was saying is its not false, for sure.

Parth (parthkohli):

I guess it is. If it's not true, then it is false.

hartnn (hartnn):

why not ? there can be incomplete information, and that info. can be true.

OpenStudy (lastdaywork):

The equation (2x-5)/(x-4) = 3 is linear. Hence, your method is a valid proof. But if the equation wasn't linear, then your solution would have been called "incomplete"

Parth (parthkohli):

I mean a counterexample to that statement is \(x = 7\). If there is a counterexample, then the statement is false.

OpenStudy (anonymous):

@LastDayWork perhaps you meant one-to-one? (2x-5)/(x-4) certainly isn't linear

hartnn (hartnn):

its not counter example if x^2 =1, x =1 true if x^2 = 1, x= -1 true if x^2=1, x =1 or x=-1 true if x^2 =1 , x=0 false

OpenStudy (nincompoop):

@ParthKohli both are true in your example you did not set a parameter that both x's be satisfied

OpenStudy (lastdaywork):

I mean it can be reduced to linear, or one-one..

hartnn (hartnn):

(2x-5)/(x-4) = 3 is not linear (2x-5)=3(x-4) is linear.

OpenStudy (lastdaywork):

^^

Parth (parthkohli):

OK -- let me rephrase that If x equals three, then x^2 - 10x + 21 must equal 0. and If x^2 - 10x + 21 equals zero, then x must equal three.

hartnn (hartnn):

lol true again x = 3 ONLY ----> False

Parth (parthkohli):

No... look... "x must equal 3" is wrong. x can equal 7.

OpenStudy (nincompoop):

but you forgot to set iff and only iff

hartnn (hartnn):

x can =7 does not mean x must not be 3.

OpenStudy (nincompoop):

wholeheartedly agrees with @hartnn the only way it will be a false statement is if you strictly made it only x=3 to satisfy the equation

hartnn (hartnn):

^^ x= 3 only is false, as already told.

Parth (parthkohli):

By "must equal three", I did mean "three ONLY"

OpenStudy (lastdaywork):

The general statement; Suppose x = a ; then f(x) = 0 Therefore, if f(x) = 0 then x = a is wrong as there can be a number b such that; f(b) = 0 but b≠a I guess, that's what @ParthKohli implies..

hartnn (hartnn):

you should have specified 'only', else it changes things.

Parth (parthkohli):

"MUST equal three"

hartnn (hartnn):

Suppose x = a ; then f(x) = 0 Therefore, if f(x) = 0 then x = a TRUE

OpenStudy (lastdaywork):

Er.. P => Q is valid only iff whenever P is true, Q is also true

OpenStudy (anonymous):

The problem is like @ParthKohli already said. Proving the converse doesn't prove the original statement.

Parth (parthkohli):

"MUST" is the same as "only", more or less. It's a word that restricts.

OpenStudy (anonymous):

x^2−10x+21= 0 is a true statement if the conclusion is x = 3 or x = 7 x2−10x+21=0, x=3 False x2−10x+21=0, x=7 false x2−10x+21=0, x=3 or x = 7 True

OpenStudy (nincompoop):

all fat cows are female therefore, all females are fat cows a ⇒ b b ⇏ a iff you're an egomaniacal douchebag a ⇔ b

OpenStudy (anonymous):

The contrapositive is what you would need to use. In this case if \(x\neq 7\) then \((2x5)/(x-4)\neq -3\)

OpenStudy (lastdaywork):

Or, in terms of Venn diagram solution set for P should be completely inside the solution set for Q. Hence, Suppose x = a ; then f(x) = 0 Therefore, if f(x) = 0 then x = a is False

OpenStudy (anonymous):

thank you for everyone's time =]